Dean Halverson recently read that full-time college students study 20 hours each week. She decides to do a study at her university to see if there is evidence to show that this is not true at her university. A random sample of 34 students were asked to keep a diary of their activities over a period of several weeks. It was found that the average number of hours that the 34 students studied each week was 17.9 hours. The sample standard deviation of 4.6 hours. Find the p -value. The p -value should be rounded to 4-decimal places.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 20

Ha : \mu \neq 20

Test statistic (t)

= (\bar x - \mu) / s / \sqrtn

= (17.9 - 20) /4.6 / \sqrt34

Test statistic = - 2.662

2 * P(z < -2.662) = 2 * 0.0039

P-value = 0.0078‬